OpenStudy (anonymous):

Shelly sews a blanket that has an area of 170 square feet. It has 30 square blocks, each the same size. What is the approximate length of each side of a block?

OpenStudy (anonymous):

let the length of side of each square block = x ft then area of one square block = x ft * x ft = x^2 ft^2 area of 30 such square blocks = 30 * x^2 ft^2 Now the total of all 30 blocks should be equal to the area of the blanket so 30 * x^2 = 170 solve this for x and you get your answer....